Given the forecast demand and actual demand for 10-foot fishing boats and 15-foot fishing boats, determine which forecast is more accurate 10-foot fishing boats Year Forecast Demand Actual Demand YEAR 1 78 71 YEAR 2 75 80 YEAR 3 83 101 YEAR 4 84 84 YEAR 5 88 68 YEAR 6 85 73 15-foot fishing boats Year Forecast Demand Actual Demand YEAR1 78 72 YEAR2 75 79 YEAR 3 93 101 YEAR 4 80 84 YEAR 5 88 75 YEAR6 85 78 determine which forecast is more accurate
Answer #1

We will use MAD to determine which forecast is more accurate. The forecast with lesser value of MAD will be the more accurate one.

MAD stands for Mean Absolute Deviation.

Absolute deviation for a particular year is the absolute difference between Actual demand and Forecast demand.

MAD will be Sum of Absolute deviations / Number of observations.

In this case , number of observations will be 6 since number of months will be 6.

Following table presents calculated values of Absolute deviation for both cases.

Year

Forecast Demand

Actual demand

Absolute deviation

Forecast demand

Actual demand

Absolute deviation

1

78

71

7

78

72

6

2

75

80

5

75

79

4

3

83

101

18

93

101

8

4

84

84

0

80

84

4

5

88

68

20

88

75

13

6

85

73

12

85

78

7

SUM =

62

42

For the first case:

Sum of absolute deviations = 62

Therefore, MAD = 62/6 = 10.33

For the second case :

Sum of absolute deviation = 42

Therefore , MAD = 42/6 =7

Since MAD for second forecast is less, it is the more accurate one.